OpenStudy (anonymous):

use the continuity of the exponential function to show that lim x->0 of (1+tx)^(1/x) = e^t use the continuity of the exponential function to show that lim x->0 of (1+tx)^(1/x) = e^t @Mathematics

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}(1+tx)^{1/x} = e^t\]

OpenStudy (anonymous):

I don't know how to use the continuity of the exponential function to show this

OpenStudy (anonymous):

i guess i don't either. the limit is easy enough, but not sure what the question wants

OpenStudy (anonymous):

how can I show that the limit equals e^t? can I do it without using the definition of e?

OpenStudy (anonymous):

without the definition of e it doesn't even make sense. i would say \[(1+xt)^{\frac{1}{x}}=e^{\frac{1}{x}\ln(1+xt)}\] and then take the limit in the exponent. maybe that is what the question is asking for, because that does in fact use "the continuity of the exponential function"

OpenStudy (anonymous):

show that \[\lim_{x\rightarrow 0}\frac{1}{x}\ln(1+xt)=t\] and you will have your solution

OpenStudy (anonymous):

that's actually perfect! exactly what I was looking for, thank you

OpenStudy (anonymous):

oh good. yw

OpenStudy (anonymous):

and solve via l'hopital i guess